# MMformer with Adaptive Transferable Attention: Advancing Multivariate Time Series Forecasting for Environmental Applications

### v1

@@ 4.3.4 Ablation Study
To assess the contribution of each MMformer component, we conducted ablation experiments by removing the MAML and MC Dropout modules. We evaluated the models based on MSE, MAE, and MAPE using an air quality dataset.
Table 8: Ablation experiment results using a real-world air quality dataset. Results are averaged across all prediction lengths (Avg). Figures in bold are the best for that metric
@@
MMformer 0.53 0.48 26.53%
Remove MAML 0.64 0.64 36.57%
Remove MC Dropout 0.57 0.52 29.08%
Remove MAMAL and MC Dropout 1.1 0.6 33.47%
Table 8 presents the comparative performance of the original MMformer, MMformer without meta-learning, MMformer without MC Dropout, and MMformer without both meta-learning and MC Dropout. While the corresponding values of MSE, MAE, and MAPE are relatively similar, it is evident that MMformer achieves the best overall performance, with the best values of MSE, MAE, and MAPE.
Figure 27: MMformer -remove MAML Prediction in Test Dataset
Figure 28: MMformer -remove MC Dropout Prediction in Test Dataset
Figure 29: MMformer -remove MAML and MC Dropout Prediction in Test Dataset
@@
Fig.13 and Fig.27-29 offer a more detailed visualization of the predicted trajectories under the various ablation settings. A direct comparison of those figures demonstrates that excluding MAML or MC Dropout leads to more pronounced deviations from the ground truth, particularly when abrupt changes occur in the time series. Specifically, excerpts where the data fluctuate rapidly highlight lagged or oversmoothed predictions for the ablated variants, whereas the complete MMformer maintains closer alignment with observed values.
Furthermore, analysis of the five test samples that exhibit the lowest MSE (Fig.30 - 32) reinforces this observation. In these figures, even minor discrepancies become apparent as the ablated models fail to accurately capture the subtler shifts in the data, indicating a reduction in robustness compared to the original MMformer. Likewise, direct comparisons between Fig.16 and Fig.3032 (ablated versions) provide additional evidence that removing either MAML or MC Dropout compromises predictive stability and accuracy.
@@
Ablation experiment evaluation quantitative outcomes (Table 8) and visual analysis demonstrate that both MAML and MC Dropout play a critical role in enhancing the MMformer’s capability to learn complex temporal dynamics, thereby ensuring consistent and robust performance in multivariate time-series forecasting. The model’s multi-layer linear structure and MAML further help avoid overfitting. Ablation experiments, which involved removing these modules, revealed notable decreases in prediction accuracy and robustness.
@@ 5 Conclusion and Future Work
This research focuses on developing a robust, high-accuracy framework for forecasting MTS. The proposed MMformer model integrates an adaptive, transferable multi-head attention mechanism with model-agnostic meta-learning (MAML) to extract and leverage underlying temporal-spatial patterns in MTS data effectively. By adaptively fine-tuning the encoder, whose parameters are transferred to the decoder, MMformer reduces the need for extensive manual parameter tuning and ensures reliable forecasts with minimal computational overhead. We assessed the performance of MMformer in forecasting MTS related to air quality and climate. Our findings indicate that MMformer provides accurate environmental change predictions with minimal computational overhead, outperforming iTransformer, Transformer, and the traditional SARIMAX model.
Despite MMformer’s notable strengths, such as its outperformance in extended-horizon forecasting, applicability across diverse regions, and capacity to handle multiple environmental factors, it still presents certain limitations. First, although the built-in MAML mechanism is generally effective for rapid adaptation in small-scale data scenarios, ensuring robust generalization across geographically expansive regions with markedly different data distributions often necessitates training on more diverse datasets. The reason is that substantial inter-regional variation demands richer samples to learn meta-representations that exhibit broad applicability. Second, although Bayesian approximate inference helps mitigate information loss, it invariably increases the complexity of the training process and diminishes transparency for non-technical stakeholders. These shortcomings underscore the importance of further research to bolster the model’s robustness and scalability.
Future work will address these shortcomings through (1) exploring semi-supervised or transfer learning approaches to reduce data requirements and enhance the model’s regional adaptability, (2) refining the approximate inference to improve computational efficiency and interpretability. These endeavors will further strengthen MMformer’s capability to detect and address environmental concerns at both local and global scales, ultimately contributing to more targeted interventions and enhanced resilience against the adverse effects of environmental changes.
In conclusion, MMformer is a robust, precise framework for MTS analysis. Future research should concentrate on scaling the model for broader datasets and exploring its potential in other domains with complex multivariate time series dynamics.
@@
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this article.

### v2

@@ 4.3.3 Ablation Study
To assess the contribution of each component in MMformer, we conducted ablation experiments by individually removing the MAML and MC Dropout modules. The resulting models are evaluated using MSE, MAE, and MAPE metrics on the air quality dataset.
Table 11: MMformer Ablation Experiment Results Using a Real-World Air Quality Dataset. The seq_len and pred_len are set to 150 and 30 for all models. The results are rounded to three decimal places. Figures in red and bold are the best for that metric.
@@
MMformer 0.245 0.332 18.873%
Remove MAML 0.252 0.338 19.166%
Remove MC Dropout 0.235 0.338 19.555%
Remove MAMAL and MC Dropout 0.239 0.334 19.323%
Table 11 summarizes the results of the ablation study, comparing the performance of the original MMformer model with three modified versions: MMformer without meta-learning (MAML), MMformer without MC Dropout, and MMformer without both meta-learning and MC Dropout. As shown in Table 11, the original MMformer achieves the lowest MSE (0.245), MAE (0.332), and MAPE (18.873%), indicating the highest predictive accuracy among all configurations. Removing meta-learning (MAML) results in a significant decline in performance, with MSE increasing to 0.252, MAE to 0.338, and MAPE to 19.166%. Excluding MC Dropout also leads to reduced performance, though the impact is less pronounced (MSE reducing to 0.235, but MAE: 0.338, MAPE: 19.555%). The removal of both meta-learning and MC Dropout causes the MSE to decrease to 0.239, but the MAE rises to 0.334, and the MAPE to 19.323%. These findings demonstrate that meta-learning and MC Dropout are critical for improving the accuracy and robustness of the MMformer model.
Ablation studies reveal that MC Dropout negatively impacts MSE. However, its crucial role in providing uncertainty analysis for climate and environmental applications, alongside its positive effects on MAE and MAPE, collectively enhances the model’s overall performance.
Figs.18-20 offer a more detailed visualization of the predicted trajectories under the various ablation settings. A direct comparison of those figures demonstrates that excluding MAML or MC Dropout leads to more pronounced deviations from the ground truth, particularly when abrupt changes occur in the time series. Specifically, excerpts where the data fluctuates rapidly highlight lagged or oversmoothed predictions for the ablated variants, whereas the complete MMformer maintains closer alignment with observed values.
@@
Figure 20: MMformer -remove MAML and MC Dropout Prediction in Test Dataset with the 5 test samples with the lowest MSE
@@
Ablation experiment evaluation quantitative outcomes (Table 11) and visual analysis (Figs.18-20) demonstrate that both MAML and MC Dropout play a critical role in enhancing the MMformer’s capability to learn complex temporal dynamics, thereby ensuring consistent and robust performance in multivariate time-series forecasting. The model’s multi-layer linear structure and MAML further help avoid overfitting. Ablation experiments, which involved removing these modules, revealed notable decreases in prediction accuracy and robustness.
4.3.4 Robustness and Efficiency Analysis
To comprehensively evaluate MMformer’s practical applicability, we conducted a thorough analysis of its robustness and computational efficiency. Robustness is assessed by examining performance stability under varying pred_len, while efficiency is measured through computational resource consumption and time during training and inference.
A robust forecasting model should maintain high performance not only across different prediction tasks but also show low sensitivity to its internal hyperparameter settings. As shown in Table 12, we test the robustness of the model to different forecasting horizons on the Air Quality dataset with prediction lengths (pred_len) set to 30, 60, and 90. Across the same seq_len (150), d_model (768), n_heads(8), encoder_layers (6), d_ff (2048), dropout (0,2), MMformer consistently achieved the lowest MSE and MAE, demonstrating its superior stability and accuracy in handling both short-term and long-term forecasting challenges. While other models experienced more significant performance degradation as the prediction length increased, MMformer maintained a clear advantage, underscoring the robustness of its architecture in capturing temporal dependencies over extended horizons. MMformer’s low sensitivity suggests that MMformer does not require exhaustive hyperparameter tuning to achieve its strong results, confirming its robustness and ease of deployment.
Table 12: Robustness Comparison of Multivariate Time Series Forecasting Models on Air Quality Datasets. Values are presented to three decimal places, with optimal results for each metric in bold red and the second in bold blue. (First label column split into Dataset and Prediction Length)
Models MMformer iTransformer TimesNet PatchTST Transformer (Ours) (2024) (2023) (2023) (2017) Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Air Quality 30 0.245 0.332 0.822 0.529 0.807 0.528 0.862 0.556 0.770 0.524 60 0.779 0.472 0.901 0.565 0.875 0.552 0.957 0.602 0.835 0.556 90 0.844 0.491 0.977 0.594 0.944 0.587 1.016 0.626 0.898 0.577 Avg 0.623 0.432 0.900 0.563 0.875 0.556 0.945 0.595 0.834 0.552
Table 13: Efficiency Comparison of Multivariate Time Series Forecasting Models on Air Quality Datasets. Values are presented to three decimal places, with optimal results for each metric in bold red and the second in bold blue. (First label column split into Dataset and Prediction Length)
Models MMformer iTransformer TimesNet PatchTST Transformer (Ours) (2024) (2023) (2023) (2017) Metric Train Inference Train Inference Train Inference Train Inference Train Inference Air Quality 30 3796 621 3597 924 16581 421 11837 191 18567 2623 60 2859 414 5053 1610 19819 492 6769 192 20176 2878 90 2783 218 4108 1668 33677 543 6755 198 20973 2844 Avg 3146 418 4253 1401 23359 485 8454 194 19905 2782
For a model to be practical, especially for large-scale environmental monitoring, it must be computationally efficient. We compared MMformer’s efficiency against baselines by measuring training time, inference time, total parameter count, and peak memory usage. All experiments are conducted on the same hardware (Intel Core i9-13900kf CPU, 32GB RAM, NVIDIA RTX 3090 GPU) to ensure a fair comparison.
Table 13 presents a comprehensive comparison of model efficiency. The analysis highlights MMformer’s competitive training and inference speeds. For instance, in the most common forecasting setting (pred_len=30), MMformer’s training time is significantly lower than that of TimesNet and Transformer, and comparable to the highly efficient iTransformer. While PatchTST shows the fastest inference time, MMformer’s inference speed is practical for most real-world applications and substantially faster than other baselines.
In terms of model complexity, MMformer strikes an effective balance. The parameter count of MMformer is 26, which is comparable to the baseline models’ parameter count of around 20. During training, the peak memory usage of MMformer and TimesNet is measured at 21 GB. In contrast, models like iTransformer and Transformer of memory usage are 7 GB, and the memory usage of PatchTST is 5 GB. While MMformer is memory-intensive during the training phase, it is highly efficient in deployment, characterized by rapid inference speeds and a minimal memory footprint. The Tables 12 and 13 solidify MMformer’s case as a high-performance yet practical solution for real-world multivariate time series forecasting.
5 Discussion
This study introduces MMformer to address critical challenges in multi-regional MTS prediction, specifically model adaptability, cross-task generalization, and uncertainty quantification. Traditional methods and earlier deep learning models like RNNs remain inadequately addressed in these aspects. MMformer enhances predictive accuracy and complex spatiotemporal time series modeling by optimizing an Encoder-only architecture through integrated timestamp handling, dimension inversion, the ATMA mechanism, and MC Dropout. The core innovation lies in the synergistic combination of these components, which effectively tackles MTS prediction issues, including data heterogeneity, prediction uncertainty, and the specific demands of multi-regional multiple features applications, delivering a more robust and adaptive solution.
5.1 Findings Relative to Research’s Objectives and Prior Work
MMformer successfully achieved its research objectives, yielding significant results in practical applications. Central to its success is the ATMA mechanism, which enables the model to "learn to learn" via meta-learning, facilitating rapid adaptation to data distribution shifts and novel forecasting tasks. It effectively addresses challenges related to computational load and data heterogeneity across multiple regions’ time series analysis. Furthermore, the application of MC Dropout not only enhances model robustness but also provides crucial uncertainty quantification, addressing the lack of uncertainty quantification challenge, a significant advance for high-stakes environmental decision-making. Quantitative comparisons with leading models, including iTransformer, TimesNet, PatchTST, and standard Transformer, demonstrate superior performance across air quality, climate datasets, and PEMS datasets. Specifically, MMformer achieved at least a 76.598% reduction in MSE, a 41.696% reduction in MAE, and a 40.097% reduction in MAPE on the air quality and climate datasets. It also achieved SOTA results on PEMS03, PEMS04, and PEMS08. These findings validate MMformer’s approach in overcoming the common limitations of existing deep learning models on spatio-temporal environmental data, particularly in capturing complex dynamics, inter-variable dependencies, and handling noise and uncertainty.
5.2 Managerial Insights and Practical Implications for Policymakers
MMformer’s enhanced forecast accuracy and reliability have significant management implications for policymakers. Proven by air quality, climate, and PEMS predictions, the model aids decision-makers in understanding environmental trends, developing more effective pollution control strategies, and optimizing resource allocation. The uncertainty quantification provided is crucial for risk assessment and the formulation of responsive measures. For instance, identifying high-risk areas allows for targeted resource deployment, mitigating environmental risks, and improving public health outcomes.
5.3 Theoretical Insights for Researchers
MMformer introduces novel perspectives for deep learning applications across multiple regions for MTS prediction. The integration of the ATMA mechanism and MC Dropout offers a generalizable methodology, enhancing model adaptability, robustness, and uncertainty quantification across diverse datasets. This study underscores the critical importance of considering prediction uncertainty in spatiotemporal data analysis. These insights are poised to drive further advancements in deep learning models for complex, dynamic, and uncertain environments across multi-regional time series.
5.4 Limitations
While MMformer has demonstrated significant advancements in multi-regional MTS prediction, certain limitations persist. Firstly, for complex single-city datasets with sparse features, the meta-learning advantage of ATMA is less pronounced, resulting in low precision. Secondly, the integration of MC Dropout for uncertainty quantification may lead to a trade-off with predictive accuracy. Thirdly, despite ATMA’s adaptive capabilities, the model still faces challenges in adapting to highly non-stationary data distributions across multiple locations. Finally, the model exhibits a degree of ’black box’ behavior, limiting deeper interpretability. The following section will summarize these findings and outline future research directions.
@@ 6 Conclusion and Future Work
This study introduces MMformer, a model designed to address key challenges in multi-regional MTS forecasting, including accuracy, adaptability, robustness to complex data, and reliable uncertainty quantification. MMformer achieves significant performance improvements on real-world environmental datasets (air quality, climate) and general MTS data (PEMS) by integrating the ATMA mechanism and MC Dropout. The core innovation lies in the synergistic integration of these components and their effective application in environmental MTS forecasting, providing enhanced adaptability and uncertainty insights compared to individual or simpler combinations.
Comprehensive experiments and ablation studies demonstrate that MMformer achieves SOTA or near-SOTA results compared to established models (iTransformer, TimesNet, PatchTST, Transformer) across air quality, climate, and PEMS datasets. Specifically, on air quality and climate datasets, MMformer shows significant percentage reductions in MSE, MAE, and MAPE compared to baseline models (e.g., at least 68.182% MSE reduction, 36.641% MAE reduction, 35.343% MAPE reduction), showcasing superior error minimization and data fitting. PEMS dataset experiments validate MMformer’s generalizability, with best performance on multiple datasets. Ablation study results consistently highlight the critical role of the meta-learning component and MC Dropout in enhancing MMformer’s overall performance, accuracy, and robustness. Although iTransformer slightly outperforms on the MAPE metric for the PEMS07 dataset, MMformer consistently demonstrates strong predictive capabilities and robustness, especially when handling complex, heterogeneous, and uncertain environmental MTS data.
Future work will focus on addressing limitations identified in this study and expanding MMformer’s capabilities and applications: (1) Enhancing computational efficiency and scalability by developing more efficient, lightweight architectures and applying model distillation and sparse attention mechanisms for real-time deployment. (2) Improving handling of extreme events and outliers by integrating self-supervised learning and causal inference to improve the detection of extreme and rare events and improve model adaptability and generalization in boundary cases. (3) Increasing model interpretability by exploring explainable time series deep learning models, such as incorporating visual attention mechanisms and causal tracing techniques, to enhance model transparency and user trust. (4) Extending the model’s applicability by expanding MMformer’s transfer capabilities to handle cross-domain and multi-modal data, and integrating online and active learning to support dynamic, real-time decision-making.
@@
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this study.
@@ Acknowledgement
Appendix A Abbreviations
Abbreviations Full name
MTS Multivariate Time Series
DL Deep learning
ATMA Adaptive Transferable Multi-head Attention
MC Dropout Monte Carlo Dropout
MLPs Multi-Layer Perceptrons
FFN Feed-forward networks
SOTA State-of-the-art
AQI Air Quality Index
SO2 Sulfur dioxide
NO2 Nitrogen dioxide
CO Carbon monoxide
O3 Ozone
GLMs Generalized Linear Models
SARIMA Seasonal Autoregressive Integrated Moving Average
ARIMA Autoregressive Integrated Moving Average
MAML Model-Agnostic Meta-Learning
RNNs Recurrent Neural Networks
LSTM Long Short-Term Memory
GRU Gated Recurrent Unit
SVMs Support Vector Machines
MAE Mean Absolute Error
MSE Mean Squared Error
MAPE Mean Absolute Percentage Error
seq_len Sequence length
pred_len Prediction length
IQR Interquartile range